# Always-Sparse Training by Growing Connections with Guided Stochastic Exploration

### v1

@@ Abstract
The excessive computational requirements of modern artificial neural networks (ANNs) are posing limitations on the machines that can run them. Sparsification of ANNs is often motivated by time, memory and energy savings only during model inference, yielding no benefits during training. A growing body of work is now focusing on providing the benefits of model sparsification also during training. While these methods greatly improve the training efficiency, the training algorithms yielding the most accurate models still materialize the dense weights, or compute dense gradients during training. We propose an efficient, always-sparse training algorithm with excellent scaling to larger and sparser models, supported by its linear time complexity with respect to the model width during training and inference. Moreover, our guided stochastic exploration algorithm improves over the accuracy of previous sparse training methods. We evaluate our method on CIFAR-10/100 and ImageNet using ResNet, VGG, and ViT models, and compare it against a range of sparsification methods111The source code is available at: https://github.com/mikeheddes/guided-stochastic-exploration.
1 Introduction
@@
Artificial neural networks (ANNs) are currently the most prominent machine learning method because of their unparalleled performance on a broad range of applications, including computer vision (Voulodimos et al., 2018; O’Mahony et al., 2019), natural language processing (Young et al., 2018; Otter et al., 2020), and reinforcement learning (Arulkumaran et al., 2017; Schrittwieser et al., 2020), among many others (Liu et al., 2017; Wang et al., 2020; Zhou et al., 2020). To improve their representational powers, ANNs keep increasing in size (Neyshabur et al., 2019; Kaplan et al., 2020). Recent large language models, for example, have widths surpassing units and total parameter counts over 100 billion (Brown et al., 2020; Rae et al., 2021; Chowdhery et al., 2022). However, with their increasing sizes, the memory and computational requirements to train and make inferences with these models becomes a limiting factor (Hwang, 2018; Ahmed and Wahed, 2020).
@@
A large body of work has addressed problems arising from the immense model sizes (Reed, 1993; Gale et al., 2019; Blalock et al., 2020; Hoefler et al., 2021). Many studies look into sparsifying the model weights based on the observation that the weight distribution of a dense model is often concentrated around zero at the end of training, indicating that the majority of weights contribute little to the function being computed (Han et al., 2015). By utilizing sparse matrix representations, the model size and the number of floating-point operations (FLOPs) can be reduced dramatically. Moreover, previous work has found that for a fixed memory or parameter budget, larger sparse models outperform smaller dense models (Zhu and Gupta, 2017; Kalchbrenner et al., 2018).
Biological neural networks have also evolved to utilize sparsity, in the human brain only one in about 10 million possible connections is realized (Gerum et al., 2020). The sparsity of the brain is seen as an important property for learning efficiency (Watts and Strogatz, 1998; Bullmore and Sporns, 2009; Pessoa, 2014). To this end, sparse ANNs have shown better generalization properties (Frankle and Carbin, 2018; Chen et al., 2021), suggesting that sparsity is indeed an effective regularizer.
@@
Early works in ANN sparsity removed connections, a process called pruning, of a trained dense model based on the magnitude of the weights (Janowsky, 1989; Ström, 1997), resulting in a more efficient model for inference. Later works improved upon this technique (Guo et al., 2016; Dong et al., 2017; Yu et al., 2018), but they all require at least the cost of training a dense model, thus yielding no efficiency benefits during training. Consequently, the resulting sparse models are limited in size by the largest dense model that can be trained on a given machine.
@@
In light of the aforementioned limitation, the lottery ticket hypothesis (LTH), surprisingly, hypothesized that there exists a subnetwork within a dense over-parameterized model that, when trained with the same initial weights, will result in a sparse model with comparable accuracy to that of the dense model (Frankle and Carbin, 2018). However, the proposed method for finding a winning ticket within a dense model is very computationally intensive, as it requires training the dense model (typically multiple times) to obtain the subnetwork. Moreover, later work weakened the hypothesis for larger ANNs (Frankle et al., 2019). Despite this, it was still an important catalyst for new methods that aim to find the winning ticket more efficiently.
@@
Efficient methods for finding a winning ticket can be categorized as: pruning before training and dynamic sparse training (DST). The before-training methods prune connections from a randomly initialized dense model (Lee et al., 2018; Tanaka et al., 2020). This means that they still require a machine capable of representing and computing with the dense model. In contrast, the DST methods start with a randomly initialized sparse model and change the connections dynamically throughout training while maintaining the overall sparsity (Mocanu et al., 2018; Mostafa and Wang, 2019; Evci et al., 2020). In practice, DST methods achieve higher accuracy than the pruning before training methods (see Section 4.3). The first DST method, sparse evolutionary training (SET), periodically prunes the connections with the lowest weight magnitude and grows new connections uniformly at random (Mocanu et al., 2018). While subsequent works, such as RigL and Top-KAST, have improved upon the accuracy of SET, they do so by periodically computing dense gradients and materializing the dense weights, respectively (Evci et al., 2020; Jayakumar et al., 2020).
@@
Instead, we present a DST algorithm that is always sparse and thus scales with the number of non-zero weights, like SET, but obtains substantially higher accuracy, on par with, and at times even exceeding, that of Top-KAST and RigL. We improve upon the accuracy of previous methods using a simple, yet highly effective, method that we call guided stochastic exploration (GSE). In short, when growing new connections, our method first efficiently samples a subset of the inactive connections for exploration. It is then guided by growing the connections with the largest gradient magnitude from the sampled subset. The key insight is that when the size of the subset is on the same order as the number of active connections, that is, the weight sparsity and gradient sparsity are similar, then the accuracy of our method surpasses that of previous sparse training methods. Our method has a time complexity of with respect to the model width , where are the number of non-zero weights. When using the common Erdős–Rényi model initialization, the time complexity simplifies to just . We evaluate the accuracy of our method experimentally on CIFAR-10/100 and ImageNet using ResNet, VGG, and ViT models and compare it against a range of sparsification methods at 90%, 95%, and 98% sparsity, indicating the percentage of zero weights. Our method promises to enable the training of much larger and sparser models.
2 Related work
@@
A wide variety of methods have been proposed that aim to reduce the size of ANNs, such as dimensionality reduction of the model weights (Jaderberg et al., 2014; Novikov et al., 2015), and weight quantization (Gupta et al., 2015; Mishra et al., 2018). However, this section only covers model sparsification methods as they are the most related to our work. Following Wang et al. (2019), the sparsification methods are categorized as: pruning after training, pruning during training, pruning before training, and dynamic sparse training.
@@
The first pruning algorithms operated on dense trained models, they pruned the connections with the smallest weight magnitude (Janowsky, 1989; Thimm and Fiesler, 1995; Ström, 1997; Han et al., 2015). This method was later generalized to first-order (Mozer and Smolensky, 1988; Karnin, 1990; Molchanov et al., 2019b, a) and second-order (LeCun et al., 1989; Hassibi and Stork, 1992; Frantar et al., 2021) Taylor polynomials of the loss with respect to the weights. These methods can be interpreted as calculating an importance score for each connection based on how its removal will effect the loss (Guo et al., 2016; Dong et al., 2017; Yu et al., 2018).
@@
Gradual pruning increases the sparsity of a model during training till the desired sparsity is reached (Zhu and Gupta, 2017; Liu et al., 2021a). Kingma et al. (2015) introduced variational dropout which adapts the dropout rate of each unit during training, Molchanov et al. (2017) showed that pruning the units with the highest dropout rate is an effective way to sparsify a model. Louizos et al. (2018) propose a method based on the reparameterization trick that allows to directly optimize the norm, which penalizes the number of non-zero weights. Alternatively, DeepHoyer is a differentiable regularizer with the same minima as the norm (Yang et al., 2019). Lastly, Powerpropagation is a sparsity promoting regularizer which raises the weights to a constant power (Schwarz et al., 2021).
@@
The lottery ticket hypothesis (LTH) (Frankle and Carbin, 2018; Frankle et al., 2019) started a line of work that aims to find a sparse model by pruning a dense model before training (Liu et al., 2018). SNIP uses the sensitivity of each connection to the loss as the importance score of a connection (Lee et al., 2018). GraSP optimizes gradient flow to accelerate training (Wang et al., 2019), however, Lubana and Dick (2020) argue that it is better to preserve gradient flow instead. Tanaka et al. (2020) highlight a problem in the aforementioned methods: they suffer from layer collapse in high sparsity regimes, that is, during the pruning phase all the connections of a layer can be removed, making the model untrainable. They propose SynFlow, which prevents layer collapse by calculating how each connection contributes to the flow of information using a path regularizer (Neyshabur et al., 2015), similar to Lee et al. (2019).
@@
The methods in this last category, including ours, start with a randomly initialized sparse model and change the connections dynamically throughout training while maintaining the same sparsity. This involves periodically pruning a fraction of active connections and growing the same number of inactive connections. SET was the first method and used magnitude pruning and random growing of connections (Mocanu et al., 2018). DeepR assigns a fixed sign to each connection at initialization and prunes those connections whose sign would change during training (Bellec et al., 2018). DSR prunes connections with the lowest global weight magnitude and uses random growing, allowing the connections to be redistributed over the layers and used in layers where they contribute the most (Mostafa and Wang, 2019). SNFS improves upon the previous methods by using an informed growing criteria: it grows the connections with the largest gradient momentum magnitude (Dettmers and Zettlemoyer, 2019). RigL makes this more efficient by only periodically calculating the dense gradients (Evci et al., 2020). However, calculating the dense gradients is still a limitation for training very large and sparse models. Top-KAST addresses this limitation by sparsely updating the dense model using only the connections with the largest weight magnitude (Jayakumar et al., 2020). While Top-KAST uses sparse gradients, it requires maintaining and updating the dense model from which the sparse model is periodically reconstructed. Moreover, we will show in Section 4.3 that Top-KAST achieves lower accuracy at higher sparsities.
3 Guided stochastic exploration
@@
Our main contribution is an efficient dynamic sparse training algorithm which is always sparse, that is, at no point is the dense model materialized, all forward passes use sparse weights, and it exclusively calculates sparse gradients. The sparsity is maintained even when changing the connections because our method only computes gradients for a subset of the inactive connections. This subset is randomly sampled during each round of growing and pruning. The connections within the subset with the largest gradient magnitudes are then selected to grow. We call our method guided stochastic exploration (GSE), referring to the stochastic sampling process of the subset and the selection of the connections by largest gradient magnitude, which provides guidance during the exploration. Although we are not the first to present an always-sparse training method, for example, SET also achieves this, among these methods, ours obtains the highest accuracy—equal or higher than less efficient methods like RigL and Top-KAST.
@@
We will use the following notation and definitions. With we denote the set of all possible connections, which is never materialized; is the set of non-zero or active connections whose weights are optimized with gradient descent; is a subset of inactive connections; is the set of grown connections; and is the set of pruned connections. Figure 1 illustrates the relations between the various connection sets. The solid gray area becomes the active set in the next step. The set is dotted to indicate that it is never materialized. Note that we abuse the notation of the connection sets to be per layer or global interchangeably.
@@
GSE is designed to incorporate the strengths of both SET and RigL. While RigL typically outperforms SET, Liu et al. (2021b) showed that SET can outperform RigL when training is significantly extended. This is because RigL only explores a fraction of the connections compared to SET. The two baseline methods can be regarded as opposing extremes in their exploration strategy. On one end, SET conducts unbiased exploration, enabling it to be highly efficient. On the opposite end, RigL employs a greedy strategy by selecting connections which appear to get the largest weight magnitudes. The exploration dynamics of GSE are, in turn, determined by the interaction of the subset and the grow set. When the subset is much larger than the grow set, in the limit including all possible connections, then the exploration is nearly greedy since the sampling process has little influence on which connections are selected. On the other hand, when the sets are equal in size, then the grown connections are fully determined by the subset sampling process. The exploration strategy of GSE is thus a hybrid of random exploration guided by the selection of the largest gradient magnitude.
@@
In the remainder of this section, we will discuss which probability distributions were considered for sampling the subset and how to efficiently sample from them in the case of fully-connected layers. In Appendix E, we discuss how our method can be applied to convolutional layers simply by interpreting them as batched fully-connected layers. We conclude this section with our complete dynamic sparse training algorithm and the time complexity.
3.1 Efficient subset sampling
@@
For our method to be always-sparse, it is critical that we can sample a subset of the inactive connections during the grow step without explicitly representing a probability for each possible connection. Note that a connection is fully specified by the input and output units it connects. Thus, to ensure an efficient sampling process, we can sample connections at random by independently sampling from distributions over the input and output units of a layer, which induces an implicit joint probability distribution over all the connections in that layer. With this process the probability of sampling a connection is never explicitly represented, just the probability of sampling each unit is represented which requires only linear, instead of quadratic, memory and time to draw samples using the alias method (Vose, 1991).
@@
Formally, connections are sampled from the discrete probability distributions over the input and output units of a layer : and . Since this process could sample duplicate or active connections, the set difference with the active set is computed in time using hash tables, resulting in a subset of connections whose size is bounded from above by the number of samples drawn, containing only inactive connections.
@@
where is an all ones vector, and are the activations and gradients of the loss at the output units of the -th layer, respectively. Here, the batch size is denoted as and are the number of units in the -th layer. The second distribution is called GraEst, its probabilities are proportional to an unbiased estimate of the gradient magnitude, given as follows:
@@
where is a vector of random signs. To compute the probabilities, both distributions can reuse the computation of the activations and the gradients during the forward and backward pass, respectively. A more detailed discussion of the distributions is provided in Appendix A. In Section 4.2, we assess which of these probability distributions is the most appropriate for sparse training.
3.2 Dynamic sparse training

### v2

@@ Abstract
The excessive computational requirements of modern artificial neural networks (ANNs) are posing limitations on the machines that can run them.
@@
Sparsification of ANNs is often motivated by time, memory and energy savings only during model inference, yielding no benefits during training. A growing body of work is now focusing on providing the benefits of model sparsification also during training. While these methods greatly improve the training efficiency, the training algorithms yielding the most accurate models still materialize the dense weights, or compute dense gradients during training.
We propose an efficient, always-sparse training algorithm with excellent scaling to larger and sparser models, supported by its linear time complexity with respect to the model width during training and inference.
Moreover, our guided stochastic exploration algorithm improves over the accuracy of previous sparse training methods.
We evaluate our method on the CIFAR-10/100 and ImageNet classification tasks using ResNet, VGG, and ViT models, and compare it against a range of sparsification methods333The source code is available at:
https://github.com/mikeheddes/guided-stochastic-exploration.
Index Terms:
dynamic sparse training, neural network pruning, sparse neural networksI Introduction
@@
Artificial neural networks (ANNs) are currently the most prominent machine learning method due to their unparalleled performance in a wide range of applications, including computer vision [1, 2], natural language processing [3, 4], and reinforcement learning [5, 6], among many others [7, 8, 9]. To improve their representational powers, ANNs continue to increase in size [10, 11]. Recent large language models, for example, have widths exceeding units and total parameter counts over 100 billion [12, 13, 14]. However, with their increasing sizes, the memory and computational requirements to train and make inferences with these models become a limiting factor [15, 16].
@@
A large body of work has addressed problems that arise from the immense model sizes [17, 18, 19, 20]. Many studies look into sparsifying the model weights based on the observation that the weight distribution of a dense model is often concentrated around zero at the end of training, indicating that the majority of weights contribute little to the function being computed [21]. Using sparse matrix representations, the model size and the number of floating-point operations (FLOPs) can be dramatically reduced. Moreover, previous work has found that for a fixed memory or parameter budget, larger sparse models outperform smaller dense models [22, 23].
@@
Early works in ANN sparsity removed connections, a process called pruning, of a trained dense model based on the magnitude of the weights [24, 25], resulting in a more efficient model for inference. Later works improved upon this technique [26, 27, 28], but all require at least the cost of training a dense model, thus providing no efficiency benefits during training. Consequently, the resulting sparse models are limited in size by the largest dense model that can be trained on a given machine.
@@
In light of the aforementioned limitation, the lottery ticket hypothesis (LTH), surprisingly, hypothesized that there exists a subnetwork within a dense over-parameterized model that, when trained with the same initial weights, will result in a sparse model with accuracy comparable to that of the dense model [29]. However, the proposed method for finding a winning ticket within a dense model is very computationally intensive, as it requires training the dense model (typically multiple times) to obtain the subnetwork. Moreover, later work weakened the hypothesis for larger ANNs [30]. Despite this, it was still an important catalyst for new methods that aim to find the winning ticket more efficiently.
@@
Efficient methods for finding a winning ticket can be categorized as pruning before training and dynamic sparse training (DST). The before-training methods prune connections from a randomly initialized dense model [31, 32]. This means that they still require a machine capable of representing and computing with the dense model. In contrast, DST methods start with a randomly initialized sparse model and dynamically change the connections throughout training while maintaining the overall sparsity [33, 34, 35], as illustrated in Figure 1. In practice, DST methods achieve higher accuracy than pruning before training methods (see Section IV-B). The first DST method, sparse evolutionary training (SET), periodically prunes the connections with the lowest weight magnitude and grows new connections uniformly at random [33]. Subsequent work, such as RigL and Top-KAST [35, 36], has improved the accuracy of SET. However, RigL periodically computes dense gradients, and while Top-KAST always computes sparse gradients, it still materializes dense weights during training.
@@
Instead, we present a DST algorithm that is always sparse, i.e., both the gradients and the weights are sparse at all times. The training and inference cost of our method therefore scales with the number of non-zero weights, like SET, but obtains substantially higher accuracy, on par with, and at times even exceeding, that of Top-KAST and RigL. We improve the accuracy of previous methods using a simple, yet effective method that we call guided stochastic exploration (GSE). In short, when new connections are added, our method first efficiently samples a subset of inactive connections for exploration. It is then guided by growing the connections with the largest gradient magnitude from the sampled subset. The key insight is that when the size of the subset is on the same order as the number of active connections, that is, the weight sparsity and gradient sparsity are similar, then the accuracy of our method surpasses that of previous sparse training methods. Our method has a time complexity of with respect to the width of the model , where is the number of non-zero weights. When using the Erdős–Rényi sparse model initialization [33], the time complexity is reduced to only . We evaluated the classification accuracy of our method on the CIFAR-10/100 and ImageNet datasets using ResNet, VGG, and ViT models and compared it with a range of sparsification methods at 90%, 95%, and 98% sparsity, indicating the percentage of zero weights. Our method promises to enable the training of larger and sparser models.
II Related work
@@
A wide variety of methods have been proposed that aim to reduce the size of ANNs, such as dimensionality reduction of the model weights [37, 38], and weight quantization [39, 40]. However, this section specifically covers model sparsification methods as they are the most related to our work. Following Wang et al. [41], the sparsification methods are categorized as: pruning after training, pruning during training, pruning before training, and dynamic sparse training.
@@
The first pruning algorithms operated on dense trained models, they prune the connections with the smallest weight magnitude [24, 42, 25, 21]. This method was later generalized to first-order [43, 44, 45, 46] and second-order [47, 48, 49] Taylor polynomials of the loss with respect to the weights. These methods can be interpreted as calculating an importance score for each connection based on how its removal will affect the loss [26, 27, 28].
@@
Gradual pruning increases the sparsity of a model during training untill the desired sparsity is reached [22, 50]. Kingma et al. [51] introduced variational dropout which adapts the dropout rate of each unit during training, Molchanov et al. [52] showed that pruning the units with the highest dropout rate is an effective way to sparsify a model. Louizos et al. [53] propose a method based on the reparameterization trick that allows to directly optimize the norm, which penalizes the number of non-zero weights. Alternatively, DeepHoyer is a differentiable regularizer with the same minima as the norm [54]. Lastly, Powerpropagation is a sparsity-promoting regularizer that raises weights to a constant power [55].
@@
The lottery ticket hypothesis (LTH) [29, 30] started a line of work that aims to find a sparse model by pruning a dense model before training [56]. SNIP uses the sensitivity of each connection to the loss as the importance score of a connection [31]. GraSP optimizes gradient flow to accelerate training [41], however, Lubana et al. [57] argue that it is better to preserve gradient flow instead. Gradient flow refers to the norm of the gradients after pruning. Tanaka et al. [32] highlight a problem in the aforementioned methods: they suffer from layer collapse in high-sparsity regimes, that is, during the pruning phase all the connections of a layer can be removed, making the model untrainable. They propose SynFlow, which prevents layer collapse by calculating how each connection contributes to the flow of information using a path regularizer [58], similar to Lee et al. [59].
@@
The methods in this last category, including ours, start with a randomly initialized sparse model and dynamically change the connections throughout training while maintaining the same sparsity. This involves periodically pruning a fraction of active connections and growing the same number of inactive connections. This process is illustrated in Figure 1. SET was the first method and used magnitude pruning and random growing of connections [33]. DeepR assigns a fixed sign to each connection at initialization and prunes those connections whose sign would change during training [60]. DSR prunes connections with the lowest global weight magnitude and uses random growing, allowing the connections to be redistributed over the layers and used in layers where they contribute the most [34]. SNFS improves upon the previous methods by using an informed growing criteria: it grows the connections with the largest gradient momentum magnitude [61]. RigL makes this more efficient by periodically calculating the dense gradients [35]. However, calculating the dense gradients is still a limitation for training very large and sparse models. Top-KAST addresses this limitation by sparsely updating the dense model using only the connections with the largest weight magnitude [36]. Although Top-KAST uses sparse gradients, it requires maintaining and updating the dense model from which the sparse model is periodically reconstructed. Moreover, we will show in Section IV-B that Top-KAST achieves lower accuracy at higher sparsities.
III Guided stochastic exploration
@@
Our main contribution is an efficient dynamic sparse training algorithm which is always sparse, that is, at no point is the dense model materialized, all forward passes use sparse weights, and it exclusively calculates sparse gradients. The sparsity is maintained even while changing the connections because our method only computes gradients for a subset of the inactive connections. This subset is randomly sampled using an efficient sampling procedure during each round of growing and pruning. The connections within the subset with the largest gradient magnitudes are then selected to grow. We call our method guided stochastic exploration (GSE), referring to the stochastic sampling process of the subset and the selection of the connections by largest gradient magnitude, which provides guidance during the exploration. Although we are not the first to present an always-sparse training method, for example, SET also achieves this, among these methods, ours obtains the highest accuracy—equal or higher than less efficient methods like RigL and Top-KAST.
@@
We will use the following notation and definitions. With we denote the set of all possible connections, which is never materialized; is the set of active connections whose weights are optimized with gradient descent, while the inactive connections have an implicit weight of zero; is a subset of inactive connections; is the set of grown connections; and is the set of pruned connections. Figure 2 illustrates the relations between the various connection sets. Note that we abuse the notation of connection sets to be per layer or global interchangeably.
@@
GSE is designed to incorporate the strengths of both SET and RigL. While RigL typically outperforms SET, Liu et al. [62] showed that SET can outperform RigL when training is significantly extended. This is because RigL only explores a fraction of the connections compared to SET. The two baseline methods can be regarded as opposing extremes in their exploration strategy. On one end, SET conducts unbiased exploration, enabling it to be highly efficient. On the opposite end, RigL employs a greedy strategy by selecting connections which appear to get the largest weight magnitudes. The exploration dynamics of GSE are, in turn, determined by the interaction of the subset and the grow set. When the subset is much larger than the grow set, in the limit including all possible connections, then the exploration is nearly greedy since the sampling process has little influence over which connections are selected. On the other hand, when the sets are equal in size, then the grown connections are fully determined by the subset sampling process. The exploration strategy of GSE is thus a hybrid of random exploration guided by the selection of the largest gradient magnitude.
@@
In the remainder of this section, we will discuss which probability distributions were considered for sampling the subset and how to efficiently sample from them in the case of feedforward layers. Our method naturally applies to convolutional layers by interpreting them as batched feedforward layers. See Appendix -E for an extended discussion. We conclude this section with our complete dynamic sparse training algorithm and its time complexity.
III-A Efficient subset sampling
@@
For our method to be always-sparse, it is critical that we can sample a subset of the inactive connections during the grow step without explicitly representing a probability for each possible connection. Note that a connection is fully specified by the input and output units it connects. Thus, to ensure an efficient sampling process, we can sample connections at random by independently sampling from distributions over the input and output units of a layer, which induces an implicit joint probability distribution over all the connections in that layer. With this process the probability of sampling a connection is never explicitly represented, just the probability of sampling each unit is represented which requires only linear, instead of quadratic, memory and time to draw samples using the alias method [63].
@@
Formally, connections are sampled from the discrete probability distributions over the input and output units of the -th layer: and . Since this process could sample duplicate or active connections, the set difference with the active set is computed in time using hash tables, resulting in a subset of connections whose size is upper bounded by the number of samples drawn, containing only inactive connections.
@@
where is an all-ones vector, and are the activations and gradients of the loss at the output units of the -th layer, respectively. Here, the batch size is denoted as and is the number of units in the -th layer. The second distribution is called GraEst, its probabilities are proportional to an unbiased estimate of the gradient magnitude, given as follows:
@@
where is a vector of random signs. To compute the probabilities, both distributions can reuse the computation of the activations and the gradients during the forward and backward pass, respectively. A more detailed discussion of the distributions is provided in Appendix -A. In Section IV-A, we assess which of these probability distributions is the most appropriate for sparse training.
III-B Dynamic sparse training